OpenStudy (anonymous):

ok I got 0.2 or 200 mL of milk and 1liter of water. together it equals 1.2 liters. now I have to figure out how much in liters the milk represents. can someone please help? im kinda confused by this

OpenStudy (anonymous):

@precal

OpenStudy (precal):

I am pretty sure you need to convert and I don't know how to do that...sorry out of my scope unless I look up conversions

OpenStudy (anonymous):

its ok but thanks

OpenStudy (anonymous):

@ikram002p ??? can you help??

OpenStudy (ikram002p):

well u have 200 ml = 0.2 ltr thats it :o

OpenStudy (anonymous):

well no I got all together 1.2liter and I need to figure out in liters 20% of that

OpenStudy (ikram002p):

ohh ic so milk in water / the whole volume of milk and water *100 %

OpenStudy (ikram002p):

so its 0.2/1.2 *100% =

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

that seems simpler than it was I guess

OpenStudy (ikram002p):

wait could u take a snap for the question? to make sure its correct ?

OpenStudy (anonymous):

Celine has a bottle that contains 20% of milk and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle. Define the variable used in the equation. (5 points) Part B: How many liters of milk are present in the bottle? Show your work. (5 points its part b I got problems with

OpenStudy (ikram002p):

so x/1+x *100% =20% solve for x , this is how u do it ^^

OpenStudy (anonymous):

ok but it is kinda confusing me with the 2 x's can you give me the steps and ill find the answer??

OpenStudy (anonymous):

Let b = volume of bottle Let m = the volume of milk Let w = the volume of water Given w= 1 lt. Given m = .2b b = m + w b = .2b + 1 .8b = 1 b = 1.25 1.25 -1 = .25 So there is .25 liters of milk in the bottle.

OpenStudy (anonymous):

thanks a lot